OpenStudy (anonymous):

HELP please medals award! What is the length of the base of a yield sign that has a height of 14in. and has an area of 84 sqaure inches?

OpenStudy (lncognlto):

Do you know how to find the area of a triangle?

OpenStudy (anonymous):

no i do not

OpenStudy (lncognlto):

Ok, so area of a triangle = 1/2 x b x h. So substituting in the values given, area = 84 and h = 14, what do we get?

OpenStudy (anonymous):

84= 1/2*b*14 ??

OpenStudy (lncognlto):

Yes, that correct! Now solving for b = ?

OpenStudy (anonymous):

how would u solve for b ?

OpenStudy (lncognlto):

By putting on one side of the equation like this: 84= 1/2*b*14 (84x2)/14 = b now b = ?

OpenStudy (anonymous):

12

OpenStudy (lncognlto):

Yep, that's correct, and thats the length of the base!

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

?

OpenStudy (lncognlto):

Hmm. Your yield signs are triangular, right? The ones in my country are, so I was just assuming yours are too.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont think we did something right

OpenStudy (anonymous):

I mean i think